OpenStudy (anonymous):

integral of: sin^4 dx

OpenStudy (anonymous):

look in back of the book for the "reduction formula" for sine

OpenStudy (anonymous):

See this is what I did, is it correct?

OpenStudy (anonymous):

pɐǝɹ oʇ pɹɐɥ ʎɹǝʌ sı sıɥʇ

OpenStudy (zarkon):

then rotate it :)

OpenStudy (anonymous):

:P Its just the way my scanner does it

OpenStudy (anonymous):

I'm at college atm, and their scanner is funky

OpenStudy (across):

That's correct.

OpenStudy (across):

By the way, nice handwriting.

OpenStudy (anonymous):

Thanks, the one thing I wasn't sure about it was this:

OpenStudy (amistre64):

\[\int sin^n(x)dx=-\frac{1}{n}sin^{n-1}(x)cos(x)+\frac{n-1}{n}\int\sin^{n-2}(x)dx\] maybe

OpenStudy (across):

No because\[(1-\cos(2x))^2=1-2\cos(2x)+\cos^2(x).\]

OpenStudy (anonymous):

Ah, I see

OpenStudy (anonymous):

Thanks guys

OpenStudy (across):

Typo on the last term above, but you're welcome.

OpenStudy (anonymous):

i think there is a mistake : sin^2 (x)= (1-cos(2x))/2

OpenStudy (across):

Isn't that what he wrote?